Find the regular convolution of the following sequences using tabular method:

π‘₯(𝑛)={3 βˆ’2 1 4} and β„Ž(𝑛)={2 5 3}

↑ ↑


1
Expert's answer
2021-05-31T06:19:05-0400

π‘₯(𝑛)={3 βˆ’2 1 4}

Range of "x(n)=-3 \eqslantless n\eqslantless0"

"\therefore x[-n]=" {-4 +2 -3 -1}

Range of "x(-n)=3 \eqslantless n\eqslantless0"

Conolution of x[n] and x[-n] "=y[x]=x[n] \text{\textcircled * x[-n]}"



The range of elements of y[n] will be "y[n] \varepsilon -3\eqslantless n \eqslantless3"

Hence elements can be calculated from -3 to 3 by summing up diagonals

y[-3]=4

y[-2]=12+-2=10

y[-1]= -8+-6+3=-11

y[0]=16+4+9+1=30

y[1]=-8+-6+3=-11

y[2]=12+-2=10

y[3]=4

y[n]={4,10,-11,30,-11,10,4}